ON THE PROFINITE TOPOLOGY ON NEGATIVELY CURVED GROUPS

RITA GITIK

Abstract. Let H and K be quasiconvex subgroups of a negatively curved LERF group G. It is shown that if H is malnormal in G then the double coset KH is closed in the profinite topology of G. In particular, this is true if G is the fundamental group of an atoroidal LERF hyperbolic 3-manifold M, and H is the fundamental group of a totally geodesic boundary component N of M.